Kelly Jones and Tami Crawford borrowed $15,750 on a 7-month, 8% note from Gem State Bank to open their business, JC’s Coffee House. The money was borrowed on June 1, 2017, and the note matures January 1, 2018.

Prepare the entry to record the receipt of the funds from the loan. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

June 1

Prepare the entry to accrue the interest on June 30. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

June 30

Assuming adjusting entries are made at the end of each month, determine the balance in the interest payable account at December 31, 2017.
Balance in interest payable account $

Prepare the entry required on January 1, 2018, when the loan is paid back. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

Jan. 1, 2018
Answer #1
Concepts and reason

Accounting: Accounting is a process of recording the transactions, classifying them in a specific manner, and is the process of summarizing and analyzing to interpret the results. It is a process of preserving the accounts.

Financial accounting is a process of preparing reports to provide all financial information to both the internal and external users of an organization. The financial statements that are prepared under the financial accounting are examined by an independent certified public accountants, at the year-end, who would express their opinion on the fairness of the reports shown by a company.

Transaction: Transaction is an act of buying or selling goods or rendering any service that is reliably measured in terms of money.

Journal entry: Journal entry is the recording of transactions in a systematic manner as they occur. Thus, it is a summary of all the transactions that has debit and credit aspects recorded chronologically.

Fundamentals

Notes payable: Notes payable are legal agreements between an issuer and a payee, where the issuer agrees to pay the payee the total amount on certain terms and conditions. The interest rate is paid on the face value of the note payable. This legal document is represented by promissory notes.

Interest expense: This is a non-operating expense, and it will be shown on the income statement. Generally, this expense is considered as an amount incurred by the company for using the other resources like borrowing loan from bank or purchasing a note and so on.

Golden rules of accounting:

Accounts
Personal
Real
Debit
The receiver
What comes in
All expenses and
losses
Credit
The giver
What goes out
All income and

Rules for debit and credit:

When assets increase, debit them; when assets decrease, credit them.

When liabilities increase, credit them; when liabilities decrease, debit them.

When stockholders’ equity increases, credit it; when stockholders’ equity decreases, debit it.

When expenses and losses increase, debit them; when expenses and losses decrease, credit them.

When incomes and gains increase, credit them; when incomes and gains decrease, debit them.

1)

Prepare the journal entry to record the receipt of the funds from the loan:

Credit
Debit
$15.750
Date Accounts title and explanation
1-Jun-17 Cash
Notes payable
(To record the receipts of fund
from the

Therefore, the cash account is debited for $15,750 and the notes payable is credited for $15,750.

2)

Prepare the journal entry to record the interest accrued:

Date Accounts title and explanation Debit Credit
30-Jun-17 Interest expense (15750x8%/12) $105
Interest payable
$105
(To reco

Therefore, the interest expense is debited and the interest payable is credited for $105.

3)

Calculate the balance in interest payable account:

Balanceininterestpayableaccrued}=Interestexpense×Numberofmonths=$105×7months=$735\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Balance in interest }}\\\\{\rm{payable accrued}}\\\end{array} \right\}{\rm{ = Interest expense}} \times {\rm{Number of months}}\\\\{\rm{ = \$ 105}} \times {\rm{7months}}\\\\{\rm{ = \$ 735}}\\\end{array}

Therefore, the balance in interest payable account is $735.

4)

Prepare the journal entry to record the loan paid:

Credit
Date Accounts title and explanation
1-Jan-18 Notes payable
Interest payable
Cash
(To record the loan paid)
Debit
$15,7

Therefore, the notes payable is debited for $15,750, interest payable is debited for $735, and the cash account is credited for $16,485.

Ans: Part 1

Credit
Debit
$15,750
Date Accounts title and explanation
1-Jun-17 Cash
Notes payable
(To record the receipts of fund
from the

Part 2

Date Accounts title and explanation Debit Credit
30-Jun-17 Interest expense
$105
Interest payable
$105
(To record the interes

Part 3

The balance in interest payable account is $735.

Part 4

Credit
Date Accounts title and explanation
1-Jan-18 Notes payable
Interest payable
Cash
(To record the loan paid)
Debit
$15,7